# Relation Also Knows: Rethinking the Recall and Editing of Factual Associations in Auto-Regressive Transformer Language Models

### v1

@@ Conclusion
We discover the over-generalizing problem for previous subject-focused knowledge editing methods, and we solve this problem by further exploring the role of relations in knowledge recall. As a result, we unveil the factual information encoded for relations in auto-regressive transformer language models, and we propose the RETS knowledge editing method based on the relation-focused interpretation. Our experiments demonstrate the effectiveness of RETS on solving the over-generalizing problem and provide the novel relation-focused perspective for future research on both the interpretation and editing of the auto-regressive transformer language models, breaking the domination of the subject-focused perspective.
@@ A. Cases for Causal Tracing on Relation Tokens
In Figure 7, we display the heatmaps of the average indirect effect of GPT2-XL and GPT-J on specific cases. All of the results exhibit the decisive position of the last relation token where the MLP outputs primarily contribute to the prediction and the most active MLP sublayers amass in either early or middle-late layers.
(a) GPT2-XL
(b) GPT-J
(c) GPT2-XL
(d) GPT-J
(e) GPT2-XL
(f) GPT-J
(g) GPT2-XL
(h) GPT-J
(i) GPT2-XL
(j) GPT-J
(k) GPT2-XL
(l) GPT-J
@@ C. Samples of Top Ranked Tokens through Vocabulary Lens
D. Causal tracing results on GPT-J and Llama-2.
Here we display the causal tracing results of GPT-J (6B) and Llama-2 (7B) in Figure 8, both showing the similar phenomenon with GPT2-XL.
(a) GPT-J
(b) GPT-J
(c) Llama-2
(d) Llama-2
@@ E. The Vocabulary Lens on GPT-J
(a) Average attributes rate
(b) Average token ranking
@@ F. Experimental Details
The existing metrics we use in our experiments are in accordance with Meng et al. 2022a. We introduce the construction of the new criterion R-Specificity here. Firstly, we need to supplement the factual prompts that share the same subject but differ in the relation for each edit. To achieve this, we categorize all the subjects of the target factual associations in COUNTERFACT into person, location, organization and product. Then, for each category, we manually creates several prompts with the subject in the category and the relations that never appear in COUNTERFACT. With the supplemented prompts, we can measure how the prediction of the raw model for unrelated factual associations is distorted by editing.
@@
All the parameter settings of the baselines are also consistent with Meng et al. 2022a. To highlight our relation-centered perspective, the parameter settings of learning rate and other optimization parameters of RETS are also consistent with Meng et al. 2022a, except that we edit in $36$-th ($18$)-th layer for GPT2-XL (GPT-J).
@@
We use subject constraints to restrict the editing to the certain subject by adding a least-square loss on the objective function, where the loss is to ensure least change on the key-value memories of several factual prompts. We simply collect the input vectors and output vectors of the selected MLP sublayer with 20,000 prompts from COUNTERFACT which is a subset of the ideal voluminous factual associations dataset that covers as much relations as possible.
@@ G. Evaluation Results on zsRE
Here we introduce some details for the RETS method, including the optimization process of $v_{*}^{p_{r}}$ and the deduction of the solution for $W^{l_{e}}$. The optimization process of $v_{*}^{p_{r}}$ is similar with ROME (Meng et al. 2022), including maximizing the probability for target object $o^{*}$ and minimizing the KL divergence of the predictions for the prompt "subject is a" to the initial model. The deduction for the solution for $W^{l_{e}}$ is similar with MEMIT (Meng et al. 2023).
I. Limitations
@@
We notice a few limitations in this work. Firstly, our design of the subject constraints for RETS is relatively simple and could affect the quality of text generations, calling for in-depth studies on how subject information is combined into the relation representation. Moreover, our work only applies to editing each factual association separately and it is essential to scale up to sequential or batched facts. Lastly, our work only focuses on the MLP sublayers for knowledge recall following previous works. The further sight into other components may provide a more complete explanation for the inference of the language models.
This paper:
Includes a conceptual outline and/or pseudocode description of AI methods introduced
(yes/partial/no/NA) yes
Clearly delineates statements that are opinions, hypothesis, and speculation from
objective facts and results (yes/no) yes
Provides well marked pedagogical references for less-familiare readers to gain background
necessary to replicate the paper (yes/no) yes
///Does this paper make theoretical contributions? (yes/no)/// |||no|||
All assumptions and restrictions are stated clearly and formally. (yes/partial/no)
All novel claims are stated formally (e.g., in theorem statements). (yes/partial/no)
Proofs of all novel claims are included. (yes/partial/no)
Proof sketches or intuitions are given for complex and/or novel results. (yes/partial/no)
Appropriate citations to theoretical tools used are given. (yes/partial/no)
All theoretical claims are demonstrated empirically to hold. (yes/partial/no/NA)
All experimental code used to eliminate or disprove claims is included. (yes/no/NA)
///Does this paper rely on one or more datasets? (yes/no)/// |||yes|||
A motivation is given for why the experiments are conducted on the selected datasets
(yes/partial/no/NA) yes
All novel datasets introduced in this paper are included in a data appendix.
(yes/partial/no/NA) yes
All novel datasets introduced in this paper will be made publicly available upon
publication of the paper with a license that allows free usage for research purposes.
(yes/partial/no/NA) yes
All datasets drawn from the existing literature (potentially including authors own
previously published work) are accompanied by appropriate citations. (yes/no/NA) yes
All datasets drawn from the existing literature (potentially including authors own
previously published work) are publicly available. (yes/partial/no/NA) yes
All datasets that are not publicly available are described in detail, with explanation
why publicly available alternatives are not scientifically satisficing.
(yes/partial/no/NA) NA
///Does this paper include computational experiments? (yes/no)/// |||yes|||
Any code required for pre-processing data is included in the appendix. (yes/partial/no). no
All source code required for conducting and analyzing the experiments is included in a
code appendix. (yes/partial/no) no
All source code required for conducting and analyzing the experiments will be made
publicly available upon publication of the paper with a license that allows free usage
for research purposes. (yes/partial/no) yes
All source code implementing new methods have comments detailing the implementation, with
references to the paper where each step comes from (yes/partial/no) partial
If an algorithm depends on randomness, then the method used for setting seeds is
described in a way sufficient to allow replication of results. (yes/partial/no/NA) NA
This paper specifies the computing infrastructure used for running experiments (hardware
and software), including GPU/CPU models; amount of memory; operating system; names and
versions of relevant software libraries and frameworks. (yes/partial/no) partial
This paper formally describes evaluation metrics used and explains the motivation for
choosing these metrics. (yes/partial/no) yes
This paper states the number of algorithm runs used to compute each reported result.
(yes/no) yes
Analysis of experiments goes beyond single-dimensional summaries of performance (e.g.,
average; median) to include measures of variation, confidence, or other distributional
information. (yes/no) no
The significance of any improvement or decrease in performance is judged using
appropriate statistical tests (e.g., Wilcoxon signed-rank). (yes/partial/no) no
This paper lists all final (hyper-)parameters used for each model/algorithm in the
papers experiments. (yes/partial/no/NA) yes
This paper states the number and range of values tried per (hyper-) parameter during
development of the paper, along with the criterion used for selecting the final parameter
setting. (yes/partial/no/NA) yes

### v2

@@ Conclusion
We discover the over-generalizing problem for previous subject-focused knowledge editing methods, and we solve this problem by further exploring the role of relations in knowledge recall. As a result, we unveil the factual information encoded for relations in auto-regressive transformer language models, and we propose the RETS single knowledge editing method based on the relation-focused interpretation. Our experiments demonstrate the effectiveness of RETS on solving the over-generalizing problem and provide the novel relation-focused perspective for future research on both the interpretation and editing of the auto-regressive transformer language models, breaking the domination of the subject-focused perspective.
Ethical Statement
The goal of our work is to investigate and renew the outdated or mistaken knowledge decoded in transformer language models. However, we recognize inherent risks associated with potential malicious applications like injecting harmful information. Therefore, we emphasize the importance that language models be sourced exclusively from reputable and trustworthy providers and carefully use the contents generated by these models.
Acknowledgements
This work is supported by the National Natural Science Foundation of China (No. 62472419, No. 62472420).
@@ A. Cases for Causal Tracing on Relation Tokens
In Figure 7, we display the heatmaps of the average indirect effect of GPT2-XL and GPT-J on specific cases. The heatmaps show the centers of interval of 10 patched layers. All of the results exhibit the decisive position of the last relation token where the MLP outputs primarily contribute to the prediction and the most active MLP sublayers amass in either early or middle-late layers.
@@ C. Samples of Top Ranked Tokens through Vocabulary Lens
D. Causal Tracing Results on GPT-J and Llama-2
Here we display the average causal tracing results of GPT-J (6B) and Llama-2 (7B) in Figure 7, both showing the similar phenomenon with GPT2-XL.
@@ F. Experimental Details
The existing metrics we use in our experiments are in accordance with Meng et al. 2022a. The original basic metrics include the basic Efficacy accuracy Scores to measure the success rate of target editing with $\mathbb{P}[o_{1}^{*}]>\mathbb{P}[o_{1}]$ where $o_{1}$ is the original object of <$s_{1},r_{1}$>, the Generalization accuracy score for generalization on the paraphrased statement <$s_{1},r_{1}^{\prime}$> with $\mathbb{P}[o_{1}^{*}]>\mathbb{P}[o_{1}]$, the renamed Subject Specificity S-Specificity accuracy score for specificity within neighborhood subjects <$s_{3},r_{1}$> with the original object $o_{3}$ that $\mathbb{P}[o_{1}^{*}]<\mathbb{P}[o_{3}]$.
We introduce the construction of the new criterion R-Specificity here. Firstly, we need to supplement the factual prompts that share the same subject but differ in the relation for each edit. To achieve this, we categorize all the subjects of the target factual associations in COUNTERFACT into person, location, organization and product. Then, for each category, we manually creates several prompts with the subject in the category and the relations that never appear in COUNTERFACT. With the supplemented prompts, we can measure how the prediction of the raw model for unrelated factual associations is distorted by editing.
@@
All the parameter settings of the baselines are consistent with Meng et al. 2022a and Meng et al. 2022b. To highlight our relation-centered perspective, the parameter settings of learning rate and other optimization parameters of RETS are also consistent with Meng et al. 2022a, except that we edit in $36$-th ($18$)-th layer for GPT2-XL (GPT-J). Evaluation experiments are conducted only once due to computational resource constraints.
For the implementaion of PMET for single editing, we simply adopt the calculation of $v$ vectors (Eqn. 4) from PMET and edit at a single layer instead of several layers as in Li et al. 2024a.
@@
We use subject constraints to restrict the editing to the certain subject by adding a least-square loss on the objective function, where the loss is to ensure least change on the key-value memories of several factual prompts. We simply collect the input vectors and output vectors of the selected MLP sublayer with 20,000 prompts from COUNTERFACT which is a subset of the ideal voluminous factual associations dataset that covers as much relations as possible. Notice that we omit the original $C$ in Eqn. 14 for the evaluation of single editing on Llama-2 for all related methods to reduce computational overhead, which has little to no effect on results.
@@ H. Method Details
Here we introduce some details for the RETS method, including the optimization process of $v_{*}^{p_{r}}$ and the deduction of the solution for $W^{l_{e}}$. The optimization process of $v_{*}^{p_{r}}$ is similar with ROME (Meng et al. 2022a), including maximizing the probability for target object $o^{*}$ and minimizing the KL divergence of the predictions for the prompt "subject is a" to the initial model. The deduction for the solution for $W^{l_{e}}$ is similar with MEMIT (Meng et al. 2022b).
I. Comparison with MEMIT
For fuller understanding, here we compare RETS with the representative massive editing method but in the case of single editing: MEMIT (Meng et al. 2022b) that spread the update of parameters over several layers. We evaluate on 1,000 samples of the supplemented COUNTERFACT dataset for GPT2-XL. Results are shown in Table 7. It indicates that the subject-focused massive editing method MEMIT also suffers the over-generalizing problem, which shows the deficiency of simply taking subject knowledge into consideration during editing.
Table 8: Performance of single editing for MEMIT and RETS. "raw" stands for the original model before edited. "Gen" stands for Generalization. "S-S" and "R-S" stands for S-Specificity and R-Specificity respectively.
Score Efficacy Gen S-S R-S
raw 55.2 20.2 22.0 78.7 100.0
MEMIT 78.3 94.9 79.5 77.4 61.3
RETS 79.9 100.0 71.5 68.7 79.3
J. The Trade-off Issue
We discuss the trade-off issue of RETS compared to the subject-focused methods here. The performance is mainly decided by the editing layer and the editing performance. For the editing layer, the purple line in Figure 5 illustrates the Subject Specificity performance of editing at the last-relation token across layers. It indicates that the subject specificity fluctuates across layers and is even worse in some earlier layers, thus the editing layer is not a major factor for the trade-off of performance. For the editing position, we further analyze the average cosine similarities of $k$ vectors at the last-relation position (input vectors in Equation 13) among subject specificity prompts and generalization prompts respectively. Results are shown in Table 8. $k$ similarity is 0.54 for subject specificity prompts, which is relatively high and indicates little subject information in $k$ vectors; $k$ similarity is 0.39 for generalization prompts. It’s not high enough for generalizing to paraphrases, hindering the generalization ability after editing. The above analysis suggests possible disjoint strategies of relation and subject information and joint strategy of paraphrase prompts for the improvement of precise editing.
Table 9: The cosine similarities of $k$ vectors of input prompts at the last-relation position. "S-S" stands for S-Specificity and "Gen" stands for Generalization.
S-S Prompts Gen Prompts
$k$ similarity 0.54 0.39
K. Limitations
@@ I. Comparison with MEMIT
We notice a few limitations in this work. Firstly, our design of the subject constraints for RETS is relatively simple and affect the quality of text generations. Fluency and Consistency for RETS decline due to the subject constraints that disrupt the associative memories of the original $K$ and $V$ in Eqn. 13, which is verified by the ablation study where Fluency increases sharply without the subject constraints. It calls for future in-depth studies on how subject information is combined into the relation representation. Moreover, our work only applies to editing each factual association separately and it is essential to scale up to sequential or batched facts. Lastly, our work only focuses on the MLP sublayers for knowledge recall following previous works. The further sight into other components may provide a more complete explanation for the inference of the language models.